Hi I’ve came up with an equation for the story problem but I am not very confident and would like some help. Below is the story problem

Sherry is 31 years younger than her mom. The sum of their ages is 61. How old is Sherry?

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52776)

.
(1)  M = S + 31    ("Sherry is 31 years younger than her mom.")
(2)  S + M = 61.   ("The sum of their ages is 61")


From (1), substitute M = S + 31  into equation (2). You will get


S + (S+31) = 61


2S + 31 = 61


2S = 61 - 31 = 30


S = 30/2 = 15.


Answer.  Sherry is 15 years old.

Solved.

Is everything clear to you in this solution ?



Answer by MathTherapy(10551)


Hi I’ve came COME up with an equation for the story problem but I am not very confident and would like some help. Below is the story problem

Sherry is 31 years younger than her mom. The sum of their ages is 61. How old is Sherry?
Don't you think it'd make perfect sense to share the equation you came up with?